OpenStudy (anonymous):

Note: The digit sum of a number is the sum of all its digits. For example N=1123 has a digit sum of 1+1+2+3=7.

OpenStudy (anonymous):

N = \[(1+10^{0})+(1+10^{1})+(1+10^{2})....+(1+10^{345})\]

OpenStudy (anonymous):

What is the digit sum of N?

OpenStudy (anonymous):

Can u help me if you can't then I suggest you leave

OpenStudy (callisto):

\[(1+10^{0})+(1+10^{1})+(1+10^{2})....+(1+10^{345})\]\[=346 + (1+10+100+...+ x)\]Where x = a number start with 1 followed by 345 0's. \[= 346 + 111...1\]The second term is a number with 346 1's.

OpenStudy (callisto):

It shouldn't be difficult then.